OpenStudy (anonymous):

Write the product in standard form. ( 9 + 5i)( 9 + 8i)

OpenStudy (itsonlycdeee):

(9+5i)(9+8i) = 81 + 72i + 45i = 40i^2 (But remember i^2 = -1) So distribute and what do you get?

OpenStudy (anonymous):

thank u :)

OpenStudy (itsonlycdeee):

No prob! (: